OpenStudy (anonymous):

I have another math problem that I need help with and I thank those that are helping me I have been trying to get the concept of this for 5 hours and this is my first time using this, thenext one is -(z+9)+(4z+1)=-2(z+1)

OpenStudy (turingtest):

first we want to get rid of the parentheses remember that \[-(a+b)\]is like\[-1(a+b)=-a-b\]and\[+(a+b)=+1(a+b)=a+b\]so try writing the left side without parentheses

OpenStudy (anonymous):

I still need help with it I just don't understand

OpenStudy (anonymous):

^ @TuringTest is right. Also, +(4z+1) is just 4z+1. -(z+9)+(4z+1)=-2(z+1) So, with what they said, you can rewrite this as so: -1(z+9)+(4z+1)=-2(z+1) z(-1) + (9)(-1) + 4z + 1 = z(-2) + 1(-2) -z - 9 + 4z + 1 = -2z -2 ___ ~~ ___ ~~~ Now, you can collect like terms here. Combine the normal numbers with the normal numbers and the numbers with z's with numbers with z's. -z - 9 + 4z + 1 = -2z -2 3z - 8 = -2z - 2 Now, add 2z to both sides. 5z - 8 = -2 Now, add 8 to both sides. 5z = 6 Can you tell me what z is?

OpenStudy (anonymous):

I'm still trying to understand it I don't know what Z is where do you get the other -1 from after the 9

OpenStudy (anonymous):

thats not it

OpenStudy (anonymous):

When you distribute, the -1 multiplies to both the z and the 9